### v1

@@ Abstract
Referring Image Segmentation (RIS) leveraging transformers has achieved great success on the interpretation of complex visual-language tasks. However, the quadratic computation cost makes it resource-consuming in capturing long-range visual-language dependencies. Fortunately, Mamba addresses this with efficient linear complexity in processing. However, directly applying Mamba to multi-modal interactions presents challenges, primarily due to inadequate channel interactions for the effective fusion of multi-modal data. In this paper, we propose ReMamber, a novel RIS architecture that integrates the power of Mamba with a multi-modal Mamba Twister block. The Mamba Twister explicitly models image-text interaction, and fuses textual and visual features through its unique channel and spatial twisting mechanism. We achieve the state-of-the-art on three challenging benchmarks. Moreover, we conduct thorough analyses of ReMamber and discuss other fusion designs using Mamba. These provide valuable perspectives for future research.
@@ 1 Introduction
Thanks to the powerful attention mechanisms, it has achieved great success for RIS by leveraging transformers to promote the exact recognition of multi-modal information. For example, existing works have designed transformer decoder [22, 4] or transformer encoder-decoder [53, 24, 49] to comprehensively fuse visual and linguistic features, which have achieved great progress. Nonetheless, it is a quadratic increase in both computation and memory when applying full attention in transformers [47, 5]. This leads to the limitation for resource-intensive scenarios, for example, in capturing long-range visual-language dependencies. And this is particularly important in the context of large-size images with long textual descriptions.
@@
Fortunately, recent advances of State Space Models (SSMs) [11, 44, 7, 8] has emerged as promising architectures for solving the above issue. Specifically, Mamba [8] marks a significant advancement for efficient training and inference with linear complexity, which has been incorporated into various visual tasks. However, current efforts primarily pay attention to single-modality settings, such as image classification [59, 33], biomedical image segmentation [35, 50, 42, 31], low-level vision [57, 14], and point cloud analysis [26, 56]. We pioneer the exploration of Mamba in multi-modal RIS setting, and identify that the prevalent modality token splicing in multi-modal transformers [24, 4, 53] is no longer effective, due to one fundamental deficiency in Mamba: there are insufficient interactions inside channels of specific or different tokens [8], which adversely affects the fusion of multi-modal information in RIS.
@@
To address this dilemma, we propose ReMamber, a novel referring segmentation architecture with Mamba twister. As shown in Fig. 1, ReMamber comprises several Mamba Twister blocks, which allow the model to be “aware” of the textual context at every spatial location. Each block consists of several visual state space (VSS) layers and a Twisting layer. Specifically, the VSS layers initially extract visual features, and the Twisting layer injects the textual information into the visual modality. The Twisting layer is structured into three critical components. (1) A vision-language interaction operation is designed to explicitly capture the fine-grained interactions between modalities. This is achieved by evaluating the similarity between visual and textual tokens then mapping them into a shared feature space, thus generating a multi-modal feature. (2) A hybrid feature cube is then created by concatenating the visual feature, multi-modal feature, and the expanded global textual feature. This ensures that each visual token receives uniform influence from both local and global contexts, preventing the overshadowing of subtle textual cues by predominant visual features. (3) Lastly, to address the inadequate interaction within channels in Mamba, a twisting mechanism is deployed. This process “twists” the feature cube channel- and spatial-wise, thereby enhancing interaction within and across modalities. It is accomplished through two consecutive SSMs scanning along channel and spatial dimensions, respectively.
@@
$\bullet$ We achieve state-of-the-arts on three challenging benchmarks. Moreover, we conduct thorough analyses of ReMamber and discuss other vision-language fusion designs using Mamba. These provide valuable perspectives for future research.
@@ 3.2 Architecture Overview
The pivotal aspect of Referring Image Segmentation (RIS) is the correspondence between image and textual input. To achieve this, we propose ReMamber. Fig. 2 shows the overview of our architecture. The basic block for ReMamber is the Mamba Twister block. It is a multi-modal feature fusion block that takes both visual and textual feature as input, and outputs the fused multi-modal feature representation. We extract the intermediate feature after each Mamba Twister block, and then feed it into a flexible decoder to generate the final segmentation mask. The decoder is task-invariant, so any downstream architecture can be applied. Below in Sec. 3.3, we will first introduce the key component: Mamba Twister block.
@@ 3.5 Twisting Layer for Multi-modal Fusion
where $\mathrm{SSM}_{\mathrm{channel}}$ refers to the Channel Scan. It is an SSM layer that treats the concatenated feature as ordered sequence through channel dimension, and presents 1-D selective scan along the channel. While $\mathrm{SSM}_{\mathrm{spatial}}$ denotes the Spatial Scan. It is an VSS layer and presents 2D selective scan along the two spatial dimension. $\mathbf{F}_{\text{out}}$ is then feed into the next layer for further process.
@@ 3.6 Discussion: Other Variants of Modality Fusion Designs
3.6.1 In-context Conditioning.
As shown in Fig. 3 (a), in this variant, we simply append the text sequence before the image feature, forming a longer sequence than originally being processed. In this way, the model is able to fuse image features with the previous text as a context. This is the most straightforward way to make the model aware of the text condition.
3.6.2 Attention-based Conditioning.
Attention mechanism is usually a strong baseline for sequence modeling. Here we utilize the cross-attention mechanism for two modality fusion, as shown in Fig. 3 (b), To be specific, we use the image feature as query, and the text feature as key and value, hoping to capture the correlation between image and text.
3.6.3 Norm Adaptation.
In this variant, we integrate text input by adapting the scale and bias after norm layer using FiLM [40]. We first pool a global representation from text sequence, and then use a linear projection to transform it into the scale and shift in layer normalization.
@@ 4.1 Datasets and Metrics
To assess the efficacy of our proposed approach, we executed experiments across three widely recognized datasets tailored for the referring image segmentation (RIS) task: RefCOCO [23], RefCOCO+ [23], and G-Ref [36, 39].
RefCOCO [23] is a prominent dataset in the RIS domain, comprising 19,994 images and 142,210 referring expressions associated with 50,000 objects derived from the MSCOCO [27] dataset through a two-player game [23]. The dataset is partitioned into 120,624 training expressions, 10,834 for validation, 5,657 for test set A, and 5,095 for test set B.
RefCOCO+ [23] enhances the challenge by excluding expressions with absolute location references, featuring 141,564 expressions for 49,856 objects across 19,992 images. Its split mirrors that of RefCOCO: 120,624 for training, 10,758 for validation, 5,726 for test A, and 5,489 for test B.
G-Ref [36, 39] enriches the dataset diversity with 104,560 referring expressions for 54,882 objects in 26,711 images, showcasing an average sentence length of 8.4 words with a heightened focus on location and appearance descriptions.
@@ 4.2 Implementation Details
Our model architecture is developed upon the foundations of Mamba [8] and VMamba [33], particularly adopting the VMamba-Base configuration for block and channel numbers following VMamba [33]. We adopt ImageNet pre-trained weights as initialization, and train the model in an end-to-end manner. We set the input image resolution to $480$ when compared with the state-of-the-art methods in Tab. 2, following [53, 46] for fair comparison. For other experiments (Tabs. 1, 3, 4 and 4), we set the image resolution to $256$ for faster training and ablation study without changing its property. We train the network for $50$ epochs using the AdamW optimizer with learning rate $\lambda=5e-5$ and cosine learning rate decay.
@@ 4.3 Evaluation on Other Variants of Modality Fusion Designs
As discussed in Sec. 3.6, besides our Mamba Twister, we also provide three other variant architectures for modality fusion, including In-context Conditioning, Attention-based Conditioning and Norm Adaptation. Here we provide a comprehensive analysis between these architectures, showing the advantages and disadvantages of each variant.
Table 1: Comparison with other modality fusion variants. “Attention” means for “Attention-based Conditioning”, “In-Context” for “In-Context Conditioning” and “Adaptation” for “Norm Adaptation”. Mamba Twister steadily outperforms other variants across all metrics and datasets, indicating its superior capability in capturing and integrating contextual information for more accurate segmentation.
@@
Adaptation 59.3 56.7 66.7 61.0 53.7 18.5 58.6 56.6 65.5 59.8 52.8 18.2
Mamba Twister 61.1 58.0 68.4 62.8 55.2 18.4 61.2 59.0 69.0 63.4 55.5 18.7
@@
Serving as a straightforward baseline, In-context Conditioning performs sub-optimal. This may because In-context Conditioning models image-text interaction in an implicit way. In our situation, the length of the image feature is much larger than the textual side, textual information may be diluted during the forward process. Besides, this operation does not differentiate the textual information between image patches, and would be insufficient for multi-modal fusion.
@@
(a) Input
(b) Cross-attention map (up) and local interaction map (down)
(c) Prediction
Figure 4: Cross-Attention map (up) and our local interaction map (down) comparison. Through both methods are able to predict target correctly, the cross-attention maps don’t show correct image-text correlation, while ours are able to capture this relationship accurately, indicating that Mamba Twister is able to gradually fusing the two modality.
@@ 4.4 Comparison with the State-of-the-Arts
Table 2: Comparison with state-of-the-art methods on RefCOCO, RefCOCO+, and G-Ref. The metric of oIoU is reported. Best results are highlighted in bold.
@@ 4.6 Visualization Results
Fig. 6 presents the visualization results of our method compared with the baseline method LAVT. Fig. 7 presents the visualization outcomes of our method alongside three other variants. Our ReMamber is capable of producing segmentation results with higher accuracy. In contrast, the other three variants occasionally encounter issues with inaccurate segmentation masks or are misled towards incorrect objects.
@@ 5 Conclusion
In this study, we introduce ReMamber, a novel architecture utilizing the Mamba framework in Referring Image Segmentation (RIS). It marks a significant advancement in multi-modal understanding. By integrating visual and textual information through innovative Mamba Twister blocks, our approach sets new benchmarks in image-text modality fusion. Achieving state-of-the-art results across multiple RIS datasets, our research highlights the potential of Mamba architecture in enhancing the scalability and performance of multi-modal tasks, offering promising directions for future exploration in the field.

### v2

@@ Abstract
Referring Image Segmentation (RIS) leveraging transformers has achieved great success on the interpretation of complex visual-language tasks. However, the quadratic computation cost makes it resource-consuming in capturing long-range visual-language dependencies. Fortunately, Mamba addresses this with efficient linear complexity in processing. However, directly applying Mamba to multi-modal interactions presents challenges, primarily due to inadequate channel interactions for the effective fusion of multi-modal data. In this paper, we propose ReMamber, a novel RIS architecture that integrates the power of Mamba with a multi-modal Mamba Twister block. The Mamba Twister explicitly models image-text interaction, and fuses textual and visual features through its unique channel and spatial twisting mechanism. We achieve competitive results on three challenging benchmarks with a simple and efficient architecture. Moreover, we conduct thorough analyses of ReMamber and discuss other fusion designs using Mamba. These provide valuable perspectives for future research. The code has been released at: https://github.com/yyh-rain-song/ReMamber.
@@ 1 Introduction
Thanks to the powerful attention mechanisms, it has achieved great success for RIS by leveraging transformers to promote the exact recognition of multi-modal information. For example, existing works have designed transformer decoder [22, 4] or transformer encoder-decoder [56, 24, 51, 36, 55, 35] to comprehensively fuse visual and linguistic features, which have achieved great progress. Nonetheless, it is a quadratic increase in both computation and memory when applying full attention in transformers [49, 5]. This leads to the limitation for resource-intensive scenarios, for example, in capturing long-range visual-language dependencies. And this is particularly important in the context of large-size images with long textual descriptions.
@@
Fortunately, recent advances of State Space Models (SSMs) [11, 46, 7, 8] has emerged as promising architectures for solving the above issue. Specifically, Mamba [8] marks a significant advancement for efficient training and inference with linear complexity, which has been incorporated into various visual tasks. However, current efforts primarily pay attention to single-modality settings, such as image classification [62, 33], biomedical image segmentation [37, 52, 44, 31], low-level vision [60, 14], and point cloud analysis [26, 59]. In this paper, we pioneer the exploration of Mamba in multi-modal RIS setting, and identify that the prevalent multi-modality token splicing method in transformers [24, 4, 56] is no longer effective. Since in Mamba, a fundamental deficiency exists: the interactions are insufficient between channels of different tokens [8], which adversely affects the fusion of multi-modal information in RIS.
@@
To address this dilemma, we propose ReMamber, a novel referring segmentation architecture with Mamba twister. As shown in Fig. 1, ReMamber comprises several Mamba Twister blocks, which allow the model to be “aware” of the textual context at every spatial location. Each block consists of several visual state space (VSS) layers and a Twisting layer. Specifically, the VSS layers initially extract visual features, and the Twisting layer injects the textual information into the visual modality. The Twisting layer is structured into three critical components. (1) A vision-language interaction operation is designed to explicitly capture the fine-grained interactions between modalities. This is achieved by evaluating the similarity between visual and textual tokens then mapping them into a shared feature space, thus generating a multi-modal feature. (2) A hybrid feature cube is then created by concatenating the visual feature, multi-modal feature, and global textual feature. This ensures that each visual token receives uniform influence from both local and global contexts, preventing the overshadowing of subtle textual cues by predominant visual features. (3) Lastly, to address the inadequate interaction within channels in Mamba, a twisting mechanism is deployed. This process “twists” the feature cube channel- and spatial-wise, thereby enhancing interaction within and across modalities. It is accomplished through two consecutive SSMs scanning along channel and spatial dimensions, respectively.
@@
$\bullet$ We achieve competitive results on three challenging benchmarks. Moreover, we conduct thorough analyses of ReMamber and discuss other vision-language fusion designs using Mamba. These provide valuable perspectives for future research.
@@ 3.2 Architecture Overview
The pivotal aspect of Referring Image Segmentation (RIS) is the correspondence between image and textual input. To achieve this, we propose ReMamber. Fig. 2 shows the overview of our architecture. The basic block for ReMamber is the Mamba Twister block. It is a multi-modal fusion block that takes both visual and textual feature as input, and outputs the fused multi-modal feature representation. We extract the intermediate feature after each Mamba Twister block, and then feed it into a flexible decoder to generate the final segmentation mask. The decoder is task-invariant, so any downstream architecture can be applied. Below in Sec. 3.3, we will first introduce the key component: Mamba Twister block.
@@ 3.5 Twisting Layer for Multi-modal Fusion
where $\mathrm{SSM}_{\mathrm{channel}}$ refers to the Channel Scan. It is an SSM layer that treats the concatenated feature as an ordered sequence through channel dimension, and presents 1-D selective scan along the channel. While $\mathrm{SSM}_{\mathrm{spatial}}$ denotes the Spatial Scan. It is an VSS layer and presents 2D selective scan along the two spatial dimension. $\mathbf{F}_{\text{out}}$ is then feed into the next layer for further process.
@@ 3.6 Discussion: Other Variants of Modality Fusion Designs
In-context Conditioning. As shown in Fig. 3 (a), in this variant, we simply append the text sequence before the image feature, forming a longer sequence than originally being processed. In this way, the model is able to fuse image features with the previous text as a context. This is the most straightforward way to allow the model aware of the text condition.
Attention-based Conditioning. Attention mechanism is usually a strong baseline for sequence modeling. Here we utilize the cross-attention mechanism for two modality fusion, as shown in Fig. 3 (b), To be specific, we use the image feature as query, and the text feature as key and value, hoping to capture the correlation between image and text.
Norm Adaptation. In this variant, we integrate the text input by adapting the scale and bias after norm layer using FiLM [42]. We first pool a global representation from text sequence, and then use a linear projection to transform it into the scale and shift in layer normalization.
@@ 4.1 Datasets and Metrics
Dataset. To assess the efficacy of our proposed approach, we executed experiments across three widely recognized datasets tailored for the referring image segmentation (RIS) task: RefCOCO, RefCOCO+ [23], and G-Ref [23, 38, 41]. RefCOCO [23] is a prominent dataset in the RIS domain, comprising 19,994 images and 142,210 referring expressions associated with 50,000 objects derived from the MSCOCO [27] dataset through a two-player game. RefCOCO+ enhances the challenge by excluding expressions with absolute location references, featuring 141,564 expressions for 49,856 objects across 19,992 images. G-Ref [38, 41] enriches the dataset diversity with 104,560 referring expressions for 54,882 objects in 26,711 images, showcasing an average sentence length of 8.4 words with a heightened focus on location and appearance descriptions.
@@ 4.2 Implementation Details
Our model architecture is developed upon the foundations of Mamba [8] and VMamba [33]. We adopt ImageNet pre-trained weights as initialization, and train the model in an end-to-end manner. We set the input image resolution to $480$ when compared with the state-of-the-art methods in Tab. 2, following [56, 48] for fair comparison. For other experiments (Tabs. 1, 3, 4 and 4), we set the image resolution to $256$ for faster training and ablation study without changing its property. For details about the training settings, please refer to our code. [^1] https://github.com/yyh-rain-song/ReMamber
@@ 4.3 Evaluation on Other Variants of Modality Fusion Designs
As discussed in Sec. 3.6, besides our Mamba Twister, we also provide three other variant architectures for modality fusion, including In-context Conditioning, Attention-based Conditioning and Norm Adaptation. To make a fair comparison, for all architectures, we initialize their common part (VMamba-based parameters) with the same checkpoint. The difference is that the Twisting layer is replaced by the corresponding fusion methods. The parameters of fusion modules are nearly equal, i.e., {103, 113, 118, 116} KB for {In-Context, Attention, Norm Adaptation, Twister}. Here we provide a comprehensive analysis between these architectures, showing the advantages and disadvantages of each variant.
Table 1: Comparison with other modality fusion variants. “Attention” means for “Attention-based Conditioning”, “In-Context” for “In-Context Conditioning”, and “Adaptation” for “Norm Adaptation”. Mamba Twister steadily outperforms other variants across all metrics and datasets, indicating its superior capability in capturing and integrating contextual information for more accurate segmentation.
@@
Adaptation 59.3 56.7 66.7 61.0 53.7 18.4 58.6 56.6 65.5 59.8 52.8 18.2
Mamba Twister 61.1 58.0 68.4 62.8 55.2 18.5 61.2 59.0 69.0 63.4 55.5 18.7
@@
Serving as a straightforward baseline, In-context Conditioning performs sub-optimal. This may because it models image-text interaction in an implicit way. In our situation, the length of the image feature is much larger than the textual side, textual information may be diluted during the forward process. Besides, this operation does not differentiate the textual information between image patches, and would be insufficient for multi-modal fusion.
@@
Figure 4: Cross-Attention map (up) and our local interaction map (down) comparison. Though both methods are able to predict target correctly, the cross-attention maps don’t show correct image-text correlation, while ours are able to capture this relationship accurately, indicating that Mamba Twister is able to gradually fusing the two modality.
@@ 4.4 Comparison with the State-of-the-Arts
Table 2: Comparison with state-of-the-art methods on RefCOCO, RefCOCO+, and G-Ref. The metric of oIoU is reported. Best results are in bold.
@@ 5 Visualization Results
Fig. 6 presents the visualization results of our method compared with the baseline method LAVT [56]. Fig. 7 presents the visualization outcomes of our method alongside three other variants. Our ReMamber is capable of producing segmentation results with higher accuracy. In contrast, the other three variants occasionally encounter issues with inaccurate segmentation masks or are misled towards incorrect objects.
@@ 6 Conclusion
In this study, we introduce ReMamber, a novel architecture utilizing the Mamba framework in Referring Image Segmentation (RIS). It marks a significant advancement in multi-modal understanding. By integrating visual and textual information through innovative Mamba Twister blocks, our approach sets new benchmarks in image-text modality fusion. Achieving competitive results across multiple RIS datasets, our research highlights the potential of Mamba architecture in enhancing the scalability and performance of multi-modal tasks, offering promising directions for future exploration in the field.
@@
Appendix
Appendix 0.A Speed Analysis
Here, we supplement two experiments w.r.t. model FPS and memory cost in Fig. 8 under different resolutions. ReMamber is consistently faster and requires fewer memory cost than LAVT, especially with large resolution (e.g., 1,024).
Figure 8: Comparison of inference FPS (left) and training GPU memory (right) between LAVT and our ReMamber.
Appendix 0.B Implementation Details
Here we provide more details about the implementation of our method, including detailed architectural structure, training settings and other baseline implementation.
0.B.1 Architecture Details
The whole ReMamber architecture consists of an encoder and a decoder. The encoder is made up by a patch-embedding layer with patch-size 4 and hidden dimension 128, followed by 4 Mamba Twister blocks. Each Twister block consists of several VSS Layers and a Twisting Layer. The VSS Layer number configuration is set as 2-2-15-2, with hidden dimension 128-256-512-1024, respectively.
For the decoder part, we provide two variants in our code implementation: convolution-based decoder (ReMamber_Conv) and Mamba-based decoder (ReMamber_Mamba). ReMamber_Conv uses a progressive upsampling architecture with 4 residual blocks, 2 convolutional layers in each. ReMamber_Mamba is similar with ReMamber_Conv, but uses VSS layers instead of convolutional layers. This variant is slightly faster.
0.B.2 Implementation for Other Three Variants
Here, we detail the implementation of other three architecture variants in our paper, i.e., In-context Conditioning, Attention-based Conditioning and Norm Adaptation.
In-context Conditioning. To enable the model to distinguish between two modalities, we add learnable positional embeddings to the image and text tokens separately at each layer before the Spatial SSM.
Attention-based Conditioning. In this variant, we also incorporate learnable positional embeddings. For the cross-attention block, we use image tokens as the query, and text tokens as the key and value.
Norm Adaptation. Norm Adaptation learns a global scale and bias. First, we use an MLP layer to pool a global vector from the text. This vector is then used to scale and bias the image tokens. An additional feed-forward layer is added after adjusting the scale and bias to maintain parameter size comparable to other variants.